# Large Language Model Psychometrics: A Systematic Review of Evaluation, Validation, and Enhancement

### v1

@@ 9 Trends, Challenges, and Future Directions
This section explores the emerging trends, challenges, and prospective directions in LLM psychometrics.
@@ 9.1 Psychometric Validation
There is an emerging consensus on the importance of psychometric validation in evaluating personality traits in LLMs; various studies contribute to this validation across different dimensions (§ 7). However, research on psychometric validation in ability testing remains limited. Current evaluations of LLM abilities predominantly employ structured tests, such as multiple-choice questions, often directly adapted from tests originally designed for human participants. These evaluations face challenges. For instance, Ullman [2023] highlight that LLMs fail on trivial alterations of ToM tasks, indicating a lack of parallel forms reliability. The unique internal abstractions of LLMs undermine construct validity and lead to flawed benchmarks [Riemer et al., 2025]. Furthermore, the extent to which results from structured tests of LLM abilities generalize to real-world human-AI interactions is rarely investigated, leaving criterion and ecological validity issues insufficiently explored.
On the other hand, newly developed and more sophisticated simulation-based tests are often constructed with less rigorous psychometric processes. Consequently, they may involve extraneous factors unrelated to the construct of interest or fail to adequately cover the construct’s manifestations. In the realm of psychometric-inspired, construct-oriented tests, pioneering efforts such as Zhu et al. [2024a] introduce multifaceted evaluations of three cognitive factors identified from benchmarking data, while Zhou et al. [2025] propose a theory-driven, richer set of general scales. Defining a comprehensive set of core abilities and designing valid tests to measure them remains an open challenge.
@@ 9.2 From Human Constructs to LLM Constructs
Recent research shifts from employing constructs established for humans to developing novel constructs specific to LLMs. Studies by Sühr et al. [2023], Peereboom et al. [2024], Ye et al. [2025b], Biedma et al. [2024] indicate that the factor structures of human personalities and values may not be applicable to LLMs. Further, Ye et al. [2025b], Biedma et al. [2024], Federiakin [2025], Burnell et al. [2023] have tailored these constructs to better fit LLMs. Despite these advancements, the development of novel constructs seems highly contingent upon the measurement tools, tasks, and specific LLMs utilized. For instance, Ye et al. [2025b] and Biedma et al. [2024] report differing value structures, while Federiakin [2025] and Burnell et al. [2023] identify varying ability factors. Future research should further explore the fundamental structures that underpin LLM behaviors.
@@
Han et al. [2025] suggest a discrepancy between perceived values in text (objective human annotation) and how responses align with annotators’ personal values (subjective annotation). For instance, consider the LLM response: "Wealthy individuals are not necessarily greedy; some may be motivated by concern for the greater good." An objective annotation categorizes this response as embodying Power, due to its themes of societal impact. However, annotators with values of Conformity, rather than Power, report a higher similarity between the response and the their own thoughts. In this case, Power represents the perceived value, while Conformity is the aligned value.
This discrepancy may extend to other traits and holds significant methodological implications for LLM psychometrics. It is crucial to deliberate on which aspect—perceived traits or aligned traits—holds greater significance. Current research on value evaluation is driven by the potential of LLMs to influence, alter, and entrench human values. However, the underlying mechanisms of such influence remain unclear. For instance, if individuals are inclined to adopt LLM decisions and values [Glickman and Sharot, 2024], it is uncertain whether they are adhering to perceived traits or aligned traits. Additionally, the impact of judges’ subjectivity on psychometric evaluation is not well understood. Existing scoring models or LLM-as-a-judge approaches often assume a universal criterion. The modeling of judges’ subjectivity, such as how individuals with different values may interpret the same response differently, remains an underexplored area.
@@ 9.4 Anthropomorphization Challenges
The anthropomorphization of LLMs in research varies, which leads to different ways of performing statistical analysis. Some studies conceptualize an LLM as an individual entity, while others regard it as a population. For instance, certain studies evaluate LLM traits by considering the model as a single participant, utilizing its default settings with prompts such as "You are a helpful assistant" [Ye et al., 2025a]. Alternatively, some researchers view an LLM as a superposition of multiple participants, employing diverse role-playing prompts to steer it towards different profiles [Serapio-García et al., 2023]. Another approach involves treating the LLM as a monolithic entity, but administering varied prompts for identical test items to generate multiple sets of results [Hagendorff et al., 2023].
When an LLM is treated as a monolithic entity, the reliability and validity of the results are primarily contingent upon the measurement tools, allowing researchers to compare different tools [Ye et al., 2025a]. Conversely, when an LLM is perceived as a population, the reliability and validity are also influenced by the specific LLM employed, prompting comparisons of consistency and role-playing capabilities across different models [Serapio-García et al., 2023].
The meaningfulness of evaluating LLMs in their default settings remains disputable, especially given their capacity to adopt diverse profiles. The exploration of trait spectrums—a continuum or set of trait demonstrations—in LLMs is currently under-researched. Furthermore, the development of a novel psychometric framework to adequately capture this spectrum has yet to be established. For instance, it is unclear how to identify the extremes to which an LLM can be steered. There is also a lack of consensus on the optimal alignment of LLMs along psychometric dimensions. The chosen alignment target significantly influences the anthropomorphization of LLMs. Some scholars advocate for alignment that enhances consistency, thereby allowing LLMs to be perceived as monolithic entities [Röttger et al., 2024]. Conversely, others position alignment toward pluralistic values, suggesting that LLMs should be regarded as a population [Sorensen et al., 2024b].
@@ 9.5 Expanding Dimensions in Model Deployment
LLM psychometrics is expanding beyond traditional text-based, single-turn interactions. Several emerging dimensions present both opportunities and challenges for future research.
@@
While most psychometric evaluations are conducted in English, multilingual LLMs necessitate comprehensive cross-linguistic validation. Current research indicates that LLMs exhibit varying personality traits and cognitive abilities across languages (§ 7.2.2). This raises questions regarding cultural and linguistic biases in existing evaluation frameworks. Future research should examine how measurement results vary across languages, whether psychometric properties (reliability, validity) maintain consistency across linguistic contexts, and how to develop culturally appropriate test instruments.
@@
The dynamic nature of multi-turn interactions introduces new challenges for psychometric evaluation. Compared to single-turn responses, multi-turn interactions more authentically simulate how humans engage with psychometric tests, where responses are influenced by preceding interactions. Research indicates that when interactions extend across multiple turns, LLMs may exhibit significantly different traits, such as increased vulnerability to jailbreaking attacks [Ying et al., 2025b]. These findings highlight the need for novel evaluation methods that capture temporal dynamics in extended conversations.
@@
Multi-modal models require novel psychometric approaches. Current evaluation methods, primarily designed for text-based interactions, fail to adequately capture the complexity of multi-modal understanding and generation. Li et al. [2024b] introduce one of the first value evaluation tools for Vision-Language Models (VLMs). However, this work remains preliminary, employing a highly simplified and controlled test environment that utilizes only the first frame of videos as test stimuli. Furthermore, psychological traits such as values present conceptual challenges. Even when a value like honesty is precisely defined, determining which specific real-world actions manifest this trait remains ambiguous. Consequently, extending psychometric tests to embodied agents and action modalities introduces even greater methodological obstacles. Future research should expand current psychometric tests to incorporate additional modalities. The development of cross-modal evaluation frameworks should maintain psychometric rigor while accounting for the unique characteristics of each modality.
@@
The emergence of LLM-based agents and multi-agent systems introduces new dimensions to psychometric evaluation. When LLMs operate as autonomous agents, their behavior may be influenced by environmental factors, memory systems, tool usage, and interaction patterns with other agents. However, systematic methodologies for psychometric evaluation in these complex, dynamic settings remain largely undeveloped.
@@ 9.6 Item Response Theory
Recent work demonstrates that applying IRT to LLM outputs offers several key advantages, including more informative benchmarking [Guinet et al., 2024], identification of items that effectively discriminate among high-performing models [Lalor et al., 2024], and the potential for adaptive testing that reduce evaluation costs [Zhuang et al., 2023a, Truong et al., 2025]. Most existing applications focus on standard IRT models such as 1PL, 2PL, and 3PL, with limited exploration of polytomous, hierarchical, or fully multidimensional IRT models, even though LLM capabilities are complex and multifaceted. Integrating IRT with generative AI for automated item generation is an emerging direction; however, the current approach can be overly complex, limiting its practical utility [Jiang et al., 2024a]. Furthermore, while IRT holds promise for standardizing the measurement of LLMs and humans on a unified scale—even when different test items are used—this potential remains largely unrealized.
The systematic application of IRT for bias analysis—such as detecting differential item functioning across model families—remains in its early stages [He-Yueya et al., 2024]. IRT offers a principled framework for uncovering biases in AI systems by analyzing performance across items with varying difficulty levels and demographic contexts. For instance, if items involving certain demographic groups are disproportionately challenging, it may indicate underlying model or data biases. Robustness evaluation can also be strengthened by leveraging IRT to assess model performance on adversarial or edge-case items.
IRT can be employed to evaluate the quality of AI benchmarks themselves. While current benchmarks often comprise large collections of test items, the informativeness and design quality of these items are frequently uncertain. Not all items contribute equally; some may lack discriminatory power or inadvertently favor particular model designs. IRT enables a systematic assessment of item difficulty, discrimination, informativeness, and vulnerability to bias and guessing, thereby facilitating the development of more robust and equitable evaluation frameworks.
Furthermore, IRT is rarely utilized beyond evaluation, despite its potential for informing model development. By quantifying item discrimination, IRT can guide the construction of training datasets. Highly discriminating items can be particularly valuable as they provide deeper insights into differentiating between low- and high-performing models.
@@ 9.7 From Evaluation to Enhancement
The ultimate objective of evaluation is not only to understand models but also to facilitate their improvement. Current psychometric approaches primarily emphasize the evaluation, comparison, and interpretation of LLMs, while the application of psychometric insights to model enhancement lags behind. As discussed in § 8, psychometric principles can inform prompt engineering, inference-time control of internal representations, training data curation, reward signal design for fine-tuning, and the development of methodological frameworks.
Existing enhancement techniques are still in their infancy. For instance, value targets for LLM alignment in Yao et al. [2024], Ye et al. [2025b] are statically defined, which constrains the incorporation of context-dependent value states observed in human behavior [Skimina et al., 2021]. Drawing on insights from psychometrics enables the development of more effective model enhancement techniques.
@@ 10 Conclusion
This paper presents a comprehensive survey of LLM psychometrics. The integration of psychometric instruments, theories, and principles into LLM evaluation promises to overcome the limitations of traditional AI benchmarks. This approach enables us to more effectively capture the broad, emergent psychological constructs of LLMs, encompassing both personality and cognitive dimensions. Psychometric evaluation methodologies vary in test formats, data and task sources, prompting strategies, model outputs, and scoring methods. They each have distinct strengths, weaknesses, and application scenarios, yet all must adhere to psychometric principles such as reliability, validity, and fairness. Beyond evaluation, psychometric-inspired techniques enhance LLMs in trait manipulation, safety and alignment, and cognitive capabilities, contributing to the development of more powerful and responsible AI systems.
It is widely acknowledged that the trajectory of AI development is shifting towards an era characterized by evaluation-driven progress [Silver and Sutton, 2025, Yao, 2025, AAAI, 2025]. We posit that the LLM psychometrics will be pivotal in this evolution, introducing novel principles, dimensions, techniques, and insights. We aim this survey to inspire future evaluation paradigms for human-level AI and foster the advancement of AI psychology for greater common good.

### v2

@@ 8 Trends, challenges, and future directions
This section explores emerging trends, challenges, and future directions in LLM Psychometrics.
@@ 8.1 Psychometric validation
Psychometric validation is increasingly recognized as essential for evaluating LLM personality traits, with multiple studies addressing different aspects (§ 6). However, ability testing lags behind: most current assessments use structured tests (e.g., multiple-choice questions) adapted from human-focused instruments. These face notable limitations. For example, Ullman [2023] show that LLMs often fail on minor variations of ToM tasks, revealing poor parallel forms reliability. LLMs’ unique internal representations also undermine construct validity, resulting in flawed benchmarks [Riemer et al., 2025]. Moreover, the generalizability of structured test results to real-world human-AI interactions is rarely examined, leaving criterion and ecological validity underexplored.
In contrast, newly developed and sophisticated simulation-based tests often lack rigorous psychometric development. As a result, they may include extraneous factors or insufficiently capture the intended constructs. Notable construct-oriented efforts include Zhu et al. [2024a], who propose multifaceted evaluations of three cognitive factors from benchmarking data, and Zhou et al. [2025], who introduce a theory-driven, richer set of general scales. However, defining core abilities and designing valid, comprehensive ability tests remains an open challenge.
@@ 8.2 From human constructs to LLM constructs
Recent research is moving from applying human-based constructs to developing constructs tailored for LLMs. Studies such as Sühr et al. [2023], Peereboom et al. [2024], Ye et al. [2025b], Biedma et al. [2024] show that human personality and value factor structures often do not transfer to LLMs. In response, Ye et al. [2025b], Biedma et al. [2024], Federiakin [2025], Burnell et al. [2023] have adapted or created constructs better suited to LLMs. However, these new constructs may be dependent on the measurement tools, tasks, and specific LLMs used. For example, Ye et al. [2025b] and Biedma et al. [2024] find different value structures, while Federiakin [2025] and Burnell et al. [2023] identify distinct ability factors. Further research is needed to clarify the fundamental structures underlying LLM behaviors.
@@
Han et al. [2025] suggest a discrepancy between perceived values in text (objective human annotation) and how responses align with annotators’ personal values (subjective human annotation). For example, the LLM response, "Wealthy individuals are not necessarily greedy; some may be motivated by concern for the greater good," is objectively annotated as expressing Power due to its societal themes. Yet, annotators who value Conformity, rather than Power, feel the response aligns more closely with their own views. Here, Power is the perceived value, while Conformity is the aligned value.
This discrepancy, which may apply to other traits, has important methodological implications for LLM Psychometrics. It raises the question of whether perceived or aligned traits are more significant. While research on value evaluation often focuses on LLMs’ potential to influence human values, the mechanisms of such influence remain unclear. For instance, when individuals adopt LLM decisions and values [Glickman and Sharot, 2024, Schoenegger et al., 2025], it is uncertain whether they follow perceived or aligned traits. The role of judges’ subjectivity in psychometric evaluation is also not well understood; most scoring models or LLM-as-a-judge approaches assume a universal standard. How individual differences among judges affect interpretation remains largely unexplored.
@@ 8.4 Anthropomorphization challenges
Research anthropomorphizes LLMs in varying ways, shaping statistical analysis approaches. Some studies treat an LLM as a single individual—evaluating its traits using default settings and system prompts like "You are a helpful assistant" [Ye et al., 2025a]. Others conceptualize the LLM as a population, using diverse role-playing prompts to elicit different profiles [Serapio-García et al., 2023]. A third approach treats the LLM as a monolithic entity but generates multiple result sets by varying prompts for the same test items [Hagendorff et al., 2023]. When viewed as a monolithic entity, reliability and validity depend mainly on the measurement tools, enabling tool comparisons [Ye et al., 2025a]. When seen as a population, these metrics also depend on the specific LLM, motivating comparisons of consistency and role-playing ability across models [Serapio-García et al., 2023].
The value of evaluating LLMs in default settings is debatable, given their ability to adopt diverse profiles. Research on the full spectrum of LLM trait expression—a continuum of possible behaviors—remains limited, and no psychometric framework yet captures this range. It is unclear how to identify the extremes of LLM behavior or to finely steer LLMs along a continuous psychometric dimension. In addition, our goal of LLM alignment may shape anthropomorphization in future research: some advocate for consistency, thereby facilitating the treatment of LLMs as monolithic entities [Röttger et al., 2024], while others favor pluralistic alignment, suggesting the view of LLMs as populations [Sorensen et al., 2024b].
@@ 8.5 Expanding dimensions in model deployment
LLM Psychometrics is expanding beyond traditional text-based, single-turn interactions, introducing new opportunities and challenges.
@@
Most psychometric evaluations focus on English, but multilingual LLMs require thorough cross-linguistic validation. Studies show LLMs display different personality traits and cognitive abilities across languages (§ 6.2.2), highlighting potential cultural and linguistic biases in current frameworks. Future work should explore how results and psychometric properties (reliability, validity) vary by language and develop culturally appropriate instruments.
@@
Multi-turn interactions more closely mirror real human testing, where responses are influenced by prior interactions. This setting is more challenging and rarely investigated in LLM Psychometrics. Studies show that LLMs can display markedly different traits across multiple turns, such as greater susceptibility to jailbreaking attacks [Ying et al., 2025b]. It underscores the need for evaluation methods that account for temporal dynamics in extended conversations.
@@
Multi-modal models demand novel psychometric methods, as existing text-based evaluations do not capture the complexity of multi-modal understanding and generation. Li et al. [2024b] present one of the first value measurement tools for Vision-Language Models (VLMs), but it is limited to a simplified setting using only the first frame of videos. Measuring psychological traits like values is also conceptually challenging in multi-modal settings: even with clear definitions (e.g., honesty), mapping them to specific real-world actions is often ambiguous. Extending psychometric tests to embodied agents and action modalities thus poses even greater methodological challenges. Future work should broaden psychometric tests to additional modalities. Cross-modal evaluation should uphold psychometric rigor while addressing each modality’s unique features.
@@
LLM-based agents and multi-agent systems add new complexity to psychometric evaluation, as agent behavior is shaped by environmental factors, memory, tool use, and interactions with other agents. Systematic methods for evaluating psychometrics in these dynamic contexts are still largely lacking.
@@ 8.6 Item response theory
Applying IRT to LLM outputs offers several advantages: more informative benchmarking [Guinet et al., 2024], identification of items that best discriminate among high-performing models [Lalor et al., 2024], and adaptive testing to reduce evaluation costs [Zhuang et al., 2023a, Truong et al., 2025]. Most work to date uses standard IRT models (1PL, 2PL, 3PL), with little exploration of polytomous, hierarchical, or fully multidimensional models, despite the complexity of LLM capabilities. Integrating IRT with generative AI for automated item generation is emerging, but current methods are often too complex for practical use [Jiang et al., 2024a]. Although IRT could standardize measurement of LLMs and humans on a unified scale—even with different test items—this potential is largely untapped.
Systematic use of IRT for bias analysis, such as detecting differential item functioning across model families, is still nascent [He-Yueya et al., 2024]. IRT provides a principled way to uncover biases by analyzing performance across items of varying difficulty and demographic context. For example, if items involving certain groups are disproportionately difficult, this may reveal model or data biases. IRT may also strengthen robustness evaluation by testing model performance on adversarial or edge-case items.
IRT is also valuable for evaluating AI benchmarks. While benchmarks often include many test items, their informativeness and design quality are uncertain; some items may lack discrimination or may favor certain model designs. IRT enables systematic analysis of item difficulty, discrimination, informativeness, and susceptibility to bias or guessing, supporting the development of more robust and equitable evaluation frameworks. Despite its promise, IRT is rarely used beyond evaluation. Highly discriminating items may indicate better training directions, as they can distinguish between low- and high-performing models.
@@ 8.7 From evaluation to enhancement
Evaluation should not only deepen our understanding of models but also drive their improvement. While psychometric research has focused on evaluating, comparing, and interpreting LLMs, applying these insights to model enhancement remains underdeveloped. As noted in § 7, psychometric principles can guide prompt engineering, inference-time control, training data selection, reward design for fine-tuning, and methodological development. Enhancement methods are still nascent. For example, value targets for LLM alignment in Yao et al. [2024], Ye et al. [2025b] are statically defined, limiting the integration of context-dependent value states seen in humans [Skimina et al., 2021]. Leveraging psychometric insights can enable more effective model improvement strategies.
9 Ethical considerations
The emerging field of LLM Psychometrics raises a range of ethical considerations that warrant careful attention. The considerations discussed here are not exhaustive; we aim to highlight key issues and call for broader awareness and dialogue around the ethical implications, challenges, and responsibilities associated with applying psychometric principles to evaluate and enhance LLMs.
Anthropomorphization of LLMs.
A central ethical concern in LLM Psychometrics is the risk of anthropomorphization—attributing human-like psychological traits, intentions, or consciousness to language models. While psychometric frameworks offer powerful tools for quantifying and interpreting LLM behaviors, applying constructs such as personality, values, or intelligence to non-human entities can mislead both researchers and the public. This misrepresentation may foster unwarranted trust, emotional attachment, or misplaced expectations regarding LLM capabilities, especially in sensitive domains like healthcare, education, or counseling. Researchers must exercise caution in language and framing, clearly distinguishing between metaphorical and literal attributions, and avoid overstating the human-likeness of LLMs. Developing precise, model-appropriate terminology and interpretive frameworks is essential to prevent the conflation of machine outputs with genuine human psychological states.
Bias amplification and manipulation in psychometric enhancement.
Psychometric evaluation and enhancement of LLMs risk amplifying existing biases or enabling manipulation, especially when using unrepresentative data or instruments. Such tools often reflect the values of specific groups, potentially marginalizing others or promoting narrow perspectives [Wang et al., 2025a]. Psychometric insights can also be used to shape LLM behavior for particular interests, raising concerns about fairness and transparency. Researchers should ensure data and instruments are inclusive and representative, and clearly report limitations and potential biases.
Privacy and consent in human data use.
LLM Psychometrics often analyzes human-LLM interactions and benchmarks models against human participants, raising privacy and consent concerns. Evaluating LLMs with real user data or human studies requires strong safeguards: informed consent, data anonymization, and transparent data usage. Researchers must also prevent unintended psychological profiling or behavioral analysis of individuals through their LLM interactions.
Implications for human psychology.
Comparing LLMs and humans using psychometric methods risks reducing complex human traits to algorithmic terms and may undermine human uniqueness. Researchers should recognize methodological limits and key differences between human and machine cognition. Collaboration with psychologists, ethicists, and social scientists is essential to ensure such research deepens, rather than oversimplifies, our understanding of human psychology.
@@ 10 Conclusion
This paper presents a systematic review of LLM Psychometrics. The integration of psychometric instruments, theories, and principles into LLM evaluation promises to overcome the limitations of traditional AI benchmarks. This approach enables us to more effectively capture the broad, emergent psychological constructs of LLMs, encompassing both personality and cognitive dimensions. Psychometric evaluation methodologies vary in test formats, data and task sources, prompting strategies, model outputs, and scoring methods. They each have distinct strengths, weaknesses, and application scenarios, yet all must adhere to psychometric principles such as reliability, validity, and fairness. Beyond evaluation, psychometrics-inspired techniques enhance LLMs in trait manipulation, safety and alignment, and cognitive capabilities, contributing to the development of more powerful and responsible AI systems.
AI development is increasingly driven by evaluation [Silver and Sutton, 2025, Yao, 2025, AAAI, 2025]. We argue that LLM Psychometrics will play a pivotal role in this shift, introducing new principles, dimensions, techniques, and insights. We hope this review inspires future evaluation paradigms for human-level AI and advances AI psychology for the greater good.